what is x when 25=5+2x
Show work

Answer by mananth(16946)

25=5+2x
add -5 to both sides
25-5=5-5+2x
20=2x
divide both sides by 2
20/2 = 2x/2
10=x
...
m.ananth@hotmail.ca